Bukit Barisan (07.01.2012 to 09.30.2012)

You are here

PDF versionPDF version
Year: 
2012
Quarter: 
July - September
Communications activities: 

No news

Unusual Events at the TEAM Site: 

No news

New Species: 

No news

Protocol Activities: 

Vegetation protocol:  During this quarter the vegetation monitoring has measured all plots (VGBBS1, 2, 3, 4, 5 and 6).  Each plots showed new trees recruitment 23, 22, 10, 19, 11 and 3.  The excel file has been checked and up loaded in TEAM portal.

Terrestrial Vertebrate protocol:  The camera traps data arra 2 (CTBBS-2) has been up loaded in desk TEAM.

Climate protocol:  All the climate data until September 2012 are up loaded on the TEAM portal

Protocol Problems: 

Terestrial Vertebrate protocol problems: Array 2 (CT-BBS-2) three camera traps (CT-BBS-2-02,12 and 30) were not working properly.  CT-BBS-2-02 and CT-BBS-2-12 rapidly flashing red and did not identify the memory card.  CT-BBS-2-30 when the camera turns on some sesor lights burning.

Climate Station: There have been some problems with the fan that we are trying to solve with the Technical Director.

Schedule Problems: 

No news

Logistical Problems: 

No news